# Quadrature Over-the-Air-Computing for Multimodal Dual-Stream Signal Processing

### v1

@@ Abstract
Index Terms:
Over-the-air computing, AirComp, multi-function computing, integrated communications and computing.
I Introduction
@@
However, despite significant research progress - including broadband channel adaptations [15], digital combining [16], and correlation handling [17] - one of the challenges of OTAC is essentially related to its computational capacity. In particular, current methods allow only a single function to be computed per resource instance (e.g., one antenna, one frequency, one symbol). This limitation arises from the standard practice of mapping data onto the real (in-phase) component of complex wireless signals, while leaving the imaginary (quadrature) component unused for data encoding and only for precoding. As a result, the system underutilizes the available complex signal space, and the effective computation rate does not fully harness its potential.
This architectural bottleneck echoes similar limitations that were overcome in other areas of wireless system design - by leveraging the inherent parallel domains of the complex field. For instance, quadrature spatial modulation (QSM) [18] extends conventional spatial modulation (SM) [19] by jointly and independently modulating the IQ components of transmitted symbols - effectively doubling the spectral efficiency and improving performance without increasing antenna count [20].
Inspired by such developments, we propose in this article quadrature over-the-air computing (Q-OTAC), a novel framework that enables the simultaneous computation of two independent nomographic functions and/or transmit data by leveraging the IQ domains of a single complex-valued transmission. This is enabled by encoding two streams of pre-processed symbols at the transmitter to each of the IQ domains, and applying a novel IQ-decoupled minimum mean square error (MMSE)-based combiner that independently estimates each target function at the AP.
@@
II System Model and Conventional OTAC
Consider a single-input multiple-output (SIMO) wireless uplink system with $K$ single-antenna EDs and a single AP equipped with $N$ receive antennas, as illustrated in Fig. 1.
Under perfect synchronization, the received signal vector $\mathbf{y}\in\mathbb{C}^{N\times 1}$ at the $N$ antennas of the AP is given by
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}p_{k}s_{k}+\mathbf{w}=\mathbf{H}\mathbf{P}\mathbf{s}+\mathbf{w}\in\mathbb{C}^{N\times 1},\vspace{-0.5ex}$ (1)
where $\mathbf{H}=[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}]\in\mathbb{C}^{N\times K}$ is the channel matrix concatenating the $K$ ED-to-AP channel vectors $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$, $\mathbf{P}=\mathrm{diag}(\mathbf{p})\in\mathbb{C}^{N\times N}$ is the diagonal matrix containing the local precoding weights $p_{k}\in\mathbb{C}$ of the EDs, collected as a vector in $\mathbf{p}\triangleq[p_{1},\ldots,p_{K}]^{\mathsf{T}}\in\mathbb{C}^{N\times 1}$, $\mathbf{s}=[s_{1},\ldots,s_{K}]^{\mathsf{T}}\in\mathbb{C}^{K\times 1}$ is the vector containing the transmit symbols $s_{k}$ of the $k$-th ED, and $\mathbf{w}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}_{w}\mathbf{I})\in\mathbb{C}^{N\times 1}$ is the additive white Gaussian noise (AWGN) vector with noise variance $\sigma^{2}_{w}\in\mathbb{R}^{+}$.
In this work, we consider the practical assumption where feedback is unavailable between the AP and the EDs, such that the transmit precoders $p_{k}$ to be fixed and known in prior [^1] When a feedback channel is available, the AP may optimize the precoding weights $p_{k}$ and distribute to the EDs. However, this requires additional signaling and is impractical in large-scale networks.. This eliminates coordination overhead and improves robustness in dense or latency-sensitive scenarios. Therefore, in the remainder of this article, we will consider $p_{k}=1\leavevmode\nobreak\ \forall k$.
Figure 1: Illustration of the conventional SIMO OTAC system.
II-A State-of-the-Art OTAC Scheme
In the conventional single-stream OTAC setting, the AP aims to compute a multivariate target function $f(d_{1},\ldots,d_{K}):\mathbb{R}^{K}\to\mathbb{R}$ of the data symbols $d_{1},\ldots,d_{K}\in\mathbb{R}$ at the $K$ EDs. The target function is assumed to be nomographic [6], which means that it can be decomposed as a post-processed sum of pre-processed symbols, $i.e.$,
$f(d_{1},\ldots,d_{K})\triangleq\psi\Bigg{(}\sum_{k=1}^{K}\underbrace{\varphi_{k}(d_{k})}_{\triangleq\;s_{k}}\Bigg{)}=\psi\Bigg{(}\sum_{k=1}^{K}s_{k}\Bigg{)}\in\mathbb{R},$ (2)
@@ II-A State-of-the-Art OTAC Scheme
For example, when the target function is a summation, the pre-/post-processing functions are simply identity functions, $i.e.$, $\varphi_{k}^{\mathrm{SUM}}(x)=x$ and $\psi^{\mathrm{SUM}}(x)=x$, such that
$\!\!f^{\mathrm{SUM}}(d_{1},\ldots,d_{K})\triangleq\psi^{\mathrm{SUM}}\Bigg{(}\sum_{k=1}^{K}\varphi_{k}^{\mathrm{SUM}}(d_{k})\Bigg{)}=\sum_{k=1}^{K}d_{k}.\!\!\!$ (3)
22footnotetext: Many real-valued multivariable functions used in practice – such as summation, average, weighted average, product, geometric mean, $p$-norms, and polynomials – are directly nomographic and thus well-suited for OTAC.
On the other hand, when the target function is a product, the pre-/post-processing functions are the logarithmic and exponentiation to an arbitrary base $b$, $i.e.$, $\varphi_{k}^{\mathrm{PROD}}(x)=\log_{b}(x)$ and $\psi^{\mathrm{PROD}}(x)=b^{\;\!x}=\mathrm{exp}_{b}(x)$, respectively, such that
$f^{\mathrm{PROD}}(d_{1},\ldots,d_{K})\triangleq\exp_{b}\!\Bigg{(}\sum_{k=1}^{K}\log_{b}\!\big{(}d_{k}\big{)}\!\Bigg{)}=\prod_{k=1}^{K}d_{k}.$ (4)
@@
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}s_{k}+\mathbf{w}=\sum_{k=1}^{K}\mathbf{h}_{k}\varphi_{k}(d_{k})+\mathbf{w}\in\mathbb{C}^{N\times 1},$ (5)
@@
Therefore, the objective of the AP is to remove the effect of the channel to yield the summation of only the pre-processed symbols, such that the post-processing function can be applied to find the target function output. Of course, each symbol can be estimated independently then recombined to yield the desired sum, but it should be stated clearly that this is not the objective of OTAC, and instead, the objective is to apply a single combiner vector $\mathbf{u}\in\mathbb{C}^{N\times 1}$ to the received signal to directly obtain the estimate of the target function, $i.e.$,
@@
In all, in hand of the channel matrix $\mathbf{H}$, the AP aims to construct an optimal combiner which minimizes the mean square error (MSE) of the estimated target function [^3] Assuming a feedback channel between the AP and the EDs, the precoding weights $p_{k}$ can also be jointly optimized with the combiner $\mathbf{u}$ in order to minimize the objective function. However, this consideration is out of scope of this article, and will be considered in an extended work., formulated as
$\mathbf{u}_{\mathrm{opt}}$ $=\operatorname*{arg\,min}_{\mathbf{u}}\,\mathbb{E}\!\left[\left|f(d_{1},\ldots,d_{K})-\psi(\mathbf{u}^{\mathsf{T}}\mathbf{y})\right|^{2}\right].\vspace{-1ex}$ (7)
While various state-of-the-art (SotA) methods exist for designing $\mathbf{u}$ – such as branch-and-bound, iterative, and message-passing schemes – we focus on MMSE-based linear combiners due to their simplicity in admitting a closed-form solution with a strong baseline performance [23, 24], which is given by
$\mathbf{u}_{\mathrm{MMSE}}=\big{(}\mathbf{H}\mathbf{H}^{\mathsf{H}}+\sigma^{2}_{w}\mathbf{I}_{N\times N}\big{)}^{\!-1}\!\cdot\mathbf{H}\cdot\mathbf{1}_{K\times 1}\in\mathbb{C}^{N\times 1},$ (8)
where $\mathbf{I}_{N\times N}$ is an identity matrix of size $N\times N$ , and $\mathbf{1}_{K\times 1}$ is an all-one vector of size $K\times 1$.
Since the combiner in eq. (8) is complex-valued, the combined output $\mathbf{u}^{\mathsf{T}}\mathbf{y}$ is also complex. However, since the original data symbols and the nomographic target functions in OTAC are real-valued, as defined in eq. (2), only the real part of the combined signal is relevant. Thus, in practice, the function estimate is obtained using the post-processing function only unto the real component of the combined output, $i.e.$,
$\tilde{f}(\mathbf{u};\mathbf{y})=\psi\!\left(\mathrm{Re}\{\mathbf{u}^{\mathsf{T}}\mathbf{y}\}\right)\in\mathbb{R}.\vspace{0.75ex}$ (9)
III Proposed Dual-Stream Quadrature OTAC
As can be seen in Section II, the conventional OTAC paradigm processes only a single real-valued function of scalar data symbols per transmission, fundamentally underutiliing the complex signal space by discarding the quadrature component, and also inherently limiting the computational capacity of each resource instance to one computing stream. To address this limitation, we propose a dual-stream Q-OTAC framework that encodes two real-valued data symbols – one in each of the IQ components of the complex transmit symbol at the EDs– with a novel combiner design which enables an independent computation of the two nomographic functions at the AP.
III-A Proposed Q-OTAC Dual-stream Transmitter Design
@@
$g(d^{\mathrm{\;\!g}}_{1},\ldots,d^{\mathrm{\;\!g}}_{K})$ $\triangleq\psi^{\mathrm{g}}\Big{(}\textstyle\sum_{k=1}^{K}\varphi^{\mathrm{g}}_{k}(d^{\mathrm{\;\!g}}_{k})\Big{)}\in\mathbb{R},\vspace{-0.5ex}$ (10b)
@@
$s_{k}\triangleq s^{\mathrm{f}}_{k}+js^{\mathrm{g}}_{k}=\varphi_{k}^{\mathrm{f}}(d_{k}^{\mathrm{\;\!f}})+j\varphi_{k}^{\mathrm{g}}(d_{k}^{\mathrm{\;\!g}})\in\mathbb{C},$ (11)
@@
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}\big{(}s_{k}^{\mathrm{f}}\!+\!js_{k}^{\mathrm{g}}\big{)}+\mathbf{w}=\mathbf{H}(\mathbf{s}^{\mathrm{f}}\!+\!j\mathbf{s}^{\mathrm{g}})+\mathbf{w}\!\in\!\mathbb{C}^{N\times 1}\vspace{-0.5ex}$ (12)
@@
Equation (12) the key novelty of the proposed Q-OTAC framework in the complex-valued transmit symbol containing two data streams in each of the IQ components, contrasting with the conventional OTAC in eq. (5) with only a real-valued transmit symbol composed of only one data stream.
Figure 2: Illustration of the proposed Q-OTAC system.
III-B Proposed Q-OTAC Dual-stream Combiner Design
@@
$\bm{y}\triangleq\begin{bmatrix}\Re\{\mathbf{y}\}\\ \Im\{\mathbf{y}\}\end{bmatrix}\!=\!\underbrace{\begin{bmatrix}\Re\{\mathbf{H}\}&-\Im\{\mathbf{H}\}\\ \Im\{\mathbf{H}\}&\Re\{\mathbf{H}\}\end{bmatrix}}_{\triangleq\bm{H}\in\mathbb{R}^{2N\times 2N}}\overbrace{\begin{bmatrix}\mathbf{s}^{\mathrm{f}}\\ \mathbf{s}^{\mathrm{g}}\\ \end{bmatrix}}^{\triangleq\bm{s}\in\mathbb{R}^{2N\times 1}}+\!\!\underbrace{\begin{bmatrix}\Re\{\mathbf{w}\}\\ \Im\{\mathbf{w}\}\end{bmatrix}}_{\triangleq\bm{w}\in\mathbb{R}^{2N\times 1}}\!\!\in\mathbb{R}^{2N\times 1},$
@@
In light of the above, the MMSE  [^4] As with conventional OTAC, other optimization methods for combiner design may be considered, however, in this work, we focus on the closed-form, low-complexity linear MMSE-based combiner as a proof of concept. problem is formulated as
@@
where the target function selector vectors are given by
$\bm{c}^{\mathrm{f}}\triangleq\begin{bmatrix}\mathbf{1}_{N\times 1}\\ \mathbf{0}_{N\times 1}\end{bmatrix}\in\mathbb{R}^{2N\times 1},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \bm{c}^{\mathrm{g}}\triangleq\begin{bmatrix}\mathbf{0}_{N\times 1}\\ \mathbf{1}_{N\times 1}\end{bmatrix}\in\mathbb{R}^{2N\times 1}.$ (17)
@@
IV Performance Analysis

### v2

@@ Abstract
Index Terms—  Quadrature, over-the-air computing, OTAC, AirComp, multifunctionality, multimodal.
1 Introduction
@@
However, despite significant research progress - including broadband channel adaptations [15], digital combining [16], and correlation handling [17] - one of the challenges of OTAC is essentially related to its computational capacity. In particular, current methods allow only a single function to be computed per resource instance (e.g., one antenna, one frequency, one symbol). This limitation arises from the standard practice of mapping data onto the in-phase component of complex wireless signals, while leaving the quadrature component unused for data encoding and only for precoding. As a result, the system underutilizes the available complex signal space, and the computation rate does not fully harness its potential.
This architectural bottleneck echoes similar limitations that were overcome in other areas of wireless system design - by leveraging the inherent parallel domains of the complex field. For instance, quadrature spatial modulation (QSM) [18] extends conventional spatial modulation (SM) [19] by jointly and independently modulating the IQ components of transmitted symbols, doubling the spectral efficiency and improving performance without increasing antenna count [20].
Inspired by such developments, we propose in this article quadrature over-the-air computing (Q-OTAC), a novel framework that enables the simultaneous computation of two independent nomographic functions and/or two data symbols by leveraging the IQ domains of a single complex-valued transmission. This is enabled by encoding two streams of pre-processed symbols at the transmitter to each of the IQ domains, and applying a novel IQ-decoupled minimum mean square error (MMSE)-based combiner that independently estimates each target function at the AP.
@@
2 System Model and Conventional OTAC
Consider a single-input multiple-output (SIMO) wireless uplink system with $K$ single-antenna EDs and a single AP equipped with $N$ receive antennas. Under perfect synchronization, the received signal vector $\mathbf{y}\in\mathbb{C}^{N\times 1}$ at the $N$ antennas of the AP is given by
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}p_{k}s_{k}+\mathbf{w}=\mathbf{H}\mathbf{P}\mathbf{s}+\mathbf{w}\in\mathbb{C}^{N\times 1},\vskip-2.15277pt$ (1)
where $\mathbf{H}=[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}]\in\mathbb{C}^{N\times K}$ is the channel matrix concatenating the $K$ ED-to-AP channel vectors $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$, $\mathbf{P}=\mathrm{diag}(\mathbf{p})\in\mathbb{C}^{K\times K}$ is the diagonal matrix containing the local precoding weights $p_{k}\in\mathbb{C}$ of the EDs, collected as a vector in $\mathbf{p}\triangleq[p_{1},\ldots,p_{K}]^{\mathsf{T}}\in\mathbb{C}^{K\times 1}$, $\mathbf{s}=[s_{1},\ldots,s_{K}]^{\mathsf{T}}\in\mathbb{C}^{K\times 1}$ is the vector containing the transmit symbols $s_{k}$ of the $k$-th ED, and $\mathbf{w}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}_{w}\mathbf{I})\in\mathbb{C}^{N\times 1}$ is the additive white Gaussian noise (AWGN) with variance $\sigma^{2}_{w}\in\mathbb{R}^{+}$.
In this work, we consider the practical assumption where feedback is unavailable between the AP and the EDs, such that the transmit precoders $p_{k}$ to be fixed and known a priori. This eliminates coordination overhead and improves robustness in dense or latency-sensitive scenarios. Therefore, in the remainder of this article, we will consider $p_{k}=1\penalty 10000\ \forall k$.
2.1 State-of-the-Art OTAC Scheme
In the conventional single-stream OTAC setting, the AP aims to compute a multivariate target function $f(d_{1},\ldots,d_{K}):\mathbb{R}^{K}\to\mathbb{R}$ of the data symbols $d_{1},\ldots,d_{K}\in\mathbb{R}$ at the $K$ EDs. The target function is assumed to be nomographic22footnotemark: 2 [6], which means that it can be decomposed as a post-processed sum of pre-processed symbols, $i.e.$,
$f(d_{1},\ldots,d_{K})\triangleq\psi\bigg(\sum_{k=1}^{K}\underbrace{\varphi_{k}(d_{k})}_{\triangleq\;s_{k}}\bigg)=\psi\bigg(\sum_{k=1}^{K}s_{k}\bigg)\in\mathbb{R},$
@@ 2.1 State-of-the-Art OTAC Scheme
For example, when the target function is a summation, the pre-/post-processing functions are simply identity functions, $i.e.$, $\varphi_{k}^{\mathrm{SUM}}(x)=x$ and $\psi^{\mathrm{SUM}}(x)=x$, such that $f^{\mathrm{SUM}}(d_{1},\ldots,d_{K})\!\triangleq\!\psi^{\mathrm{SUM}}\big(\sum_{k=1}^{K}\varphi_{k}^{\mathrm{SUM}}(d_{k})\big)\!=\!\sum_{k=1}^{K}d_{k}.$
11footnotetext: Many real-valued multivariable functions used in practice – such as summation, average, weighted average, product, geometric mean, $p$-norms, and polynomials – are directly nomographic and thus well-suited for OTAC.
On the other hand, when the target function is a product, the pre-/post-processing functions are the logarithmic and exponentiation to an arbitrary base $b$, $i.e.$, $\varphi_{k}^{\mathrm{PROD}}(x)=\log_{b}(x)$ and $\psi^{\mathrm{PROD}}(x)=b^{\;\!x}=\mathrm{exp}_{b}(x)$, respectively, such that $f^{\mathrm{PROD}}(d_{1},\ldots,d_{K})\triangleq\exp_{b}\!\big(\sum_{k=1}^{K}\log_{b}(d_{k})\big)\!=\!\prod_{k=1}^{K}d_{k}.$
@@
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}s_{k}+\mathbf{w}=\sum_{k=1}^{K}\mathbf{h}_{k}\varphi_{k}(d_{k})+\mathbf{w}\in\mathbb{C}^{N\times 1},\vskip-4.30554pt$ (3)
@@
Therefore, the objective of the AP is to remove the effect of the channel to yield the summation of only the pre-processed symbols, such that the post-processing function can be applied to find the target function output. Of course, each symbol can be estimated independently then recombined to yield the desired sum, but it should be stated clearly that this is not the objective of OTAC, and instead, the objective is to apply a single combiner vector $\mathbf{u}\in\mathbb{C}^{N\times 1}$ to the received signal to obtain the estimate of the target function, $i.e.$,
@@
In all, given the channel matrix $\mathbf{H}$, the AP aims to construct an optimal combiner which minimizes the mean square error (MSE) of the estimated target function [^2] Assuming a feedback channel between the AP and the EDs, the precoding weights $p_{k}$ can also be jointly optimized with the combiner $\mathbf{u}$ in order to minimize the objective function. However, this consideration is out of scope of this article, and will be considered in an extended work., formulated as
$\mathbf{u}_{\mathrm{opt}}$ $=\operatorname*{arg\,min}_{\mathbf{u}}\,\mathbb{E}\!\left[\left|f(d_{1},\ldots,d_{K})-\psi(\mathbf{u}^{\mathsf{T}}\mathbf{y})\right|^{2}\right].\vskip-8.61108pt$ (5)
While various state-of-the-art (SotA) methods exist for designing $\mathbf{u}$ – such as branch-and-bound, iterative, and message-passing schemes – we focus on MMSE-based linear combiners due to their simplicity in admitting a closed-form solution with a strong baseline performance [23, 24], with
$\mathbf{u}_{\mathrm{MMSE}}=\big(\mathbf{H}\mathbf{H}^{\mathsf{H}}+\sigma^{2}_{w}\mathbf{I}_{N\times N}\big)^{\!-1}\!\cdot\mathbf{H}\cdot\mathbf{1}_{K\times 1}\in\mathbb{C}^{N\times 1}.\vskip-4.30554pt$ (6)
Since the combiner in eq. (6) is complex-valued, the combined output $\mathbf{u}^{\mathsf{T}}\mathbf{y}$ is also complex. However, since the original data symbols and the nomographic target functions in OTAC are real-valued, as defined in eq. (LABEL:eq:OTACtargetfunction), only the real part of the combined signal is relevant. Thus, in practice, the function estimate is obtained using the post-processing function only unto the real component of the combined output, $i.e.$,
$\tilde{f}(\mathbf{u};\mathbf{y})=\psi\!\left(\mathrm{Re}\{\mathbf{u}^{\mathsf{T}}\mathbf{y}\}\right)\in\mathbb{R}.\vskip 3.22916pt$ (7)
3 Proposed Dual-Stream Q-OTAC
As can be seen in Section 2, the conventional OTAC paradigm processes only a single real-valued function of scalar data symbols per transmission, fundamentally underutiliing the complex signal space by discarding the quadrature component, and also inherently limiting the computational capacity of each resource instance to one computing stream.
To address this limitation, we propose a dual-stream quadrature over-the-air computing (Q-OTAC) framework that encodes two real-valued data symbols – one in each of the IQ components of the complex transmit symbol – with a novel combiner design which enables an independent computation of the two nomographic functions at the AP, which has been illustrated in Fig. 1, clearly highlighting the two streams.
3.1 Proposed Q-OTAC Dual-stream Transmitter Design
@@
$g(d^{\mathrm{\;\!g}}_{1},\ldots,d^{\mathrm{\;\!g}}_{K})$ $\triangleq\psi^{\mathrm{g}}\Big(\textstyle\sum_{k=1}^{K}\varphi^{\mathrm{g}}_{k}(d^{\mathrm{\;\!g}}_{k})\Big)\in\mathbb{R},\vskip-2.15277pt$ (8b)
@@
$s_{k}\triangleq s^{\mathrm{f}}_{k}+js^{\mathrm{g}}_{k}=\varphi_{k}^{\mathrm{f}}(d_{k}^{\mathrm{\;\!f}})+j\varphi_{k}^{\mathrm{g}}(d_{k}^{\mathrm{\;\!g}})\in\mathbb{C},\vskip-2.15277pt$ (9)
@@
$\mathbf{y}=\sum_{k=1}^{K}\mathbf{h}_{k}\big(s_{k}^{\mathrm{f}}\!+\!js_{k}^{\mathrm{g}}\big)+\mathbf{w}=\mathbf{H}(\mathbf{s}^{\mathrm{f}}\!+\!j\mathbf{s}^{\mathrm{g}})+\mathbf{w}\!\in\!\mathbb{C}^{N\times 1}\vskip-6.45831pt$ (10)
@@
Equation (10) illustrates the key novelty of the proposed Q-OTAC framework in the complex-valued transmit symbol containing two data streams in each of the IQ components, contrasting with the conventional OTAC in eq. (3) with only a real-valued transmit symbol composed of only one stream.
Fig. 1: Illustration of the proposed Q-OTAC system.
3.2 Proposed Q-OTAC Dual-stream Combiner Design
@@
$\bm{y}\triangleq\begin{bmatrix}\Re\{\mathbf{y}\}\\ \Im\{\mathbf{y}\}\end{bmatrix}\!=\!\underbrace{\begin{bmatrix}\Re\{\mathbf{H}\}&\hskip-6.45831pt-\Im\{\mathbf{H}\}\\ \Im\{\mathbf{H}\}&\Re\{\mathbf{H}\}\end{bmatrix}}_{\triangleq\bm{H}\in\mathbb{R}^{2N\times 2N}}\hskip-9.68747pt\overbrace{\begin{bmatrix}\mathbf{s}^{\mathrm{f}}\\ \mathbf{s}^{\mathrm{g}}\\ \end{bmatrix}}^{\triangleq\bm{s}\in\mathbb{R}^{2N\times 1}}\hskip-9.68747pt+\!\!\underbrace{\begin{bmatrix}\Re\{\mathbf{w}\}\\ \Im\{\mathbf{w}\}\end{bmatrix}}_{\triangleq\bm{w}\in\mathbb{R}^{2N\times 1}}\!\!\in\mathbb{R}^{2N\times 1},$
@@
Given the above, the MMSE  [^3] As with conventional OTAC, other optimization methods for combiner design may be considered, however, in this work, we focus on the closed-form, low-complexity linear MMSE-based combiner as a proof of concept of the proposed dual-stream framework. problem is formulated as
@@
where the auxiliary domain selector vectors $\bm{c}^{\mathrm{f}}\in\mathbb{R}^{2N\times 1}$ and $\bm{c}^{\mathrm{g}}\in\mathbb{R}^{2N\times 1}$ are simply given by
$\bm{c}^{\mathrm{f}}\triangleq\begin{bmatrix}\mathbf{1}_{N\times 1}\\ \mathbf{0}_{N\times 1}\end{bmatrix}\in\mathbb{R}^{2N\times 1},\penalty 10000\ \text{and}\penalty 10000\ \bm{c}^{\mathrm{g}}\triangleq\begin{bmatrix}\mathbf{0}_{N\times 1}\\ \mathbf{1}_{N\times 1}\end{bmatrix}\in\mathbb{R}^{2N\times 1}.$ (15)
@@
4 Performance Analysis
Fig. 2: The NMSE CDF for different multimodal dual-stream Q-OTAC scenarios for $N=20,K=20$ and $\mathrm{SNR}=15\mathrm{dB}$.